OpenStudy (anonymous):

excuse me, in triangle ABC, AB=5, COSB=-3/5 SINA=? one more condition, i'll use the tool bar sorry,cuz i dont know how to describe it in english

OpenStudy (anonymous):



OpenStudy (harsimran_hs4):

can you mark the sides of triangle and is cosb = -3/5?

terenzreignz (terenzreignz):

 Is this the diameter?

OpenStudy (anonymous):

no picture in the question

OpenStudy (anonymous):

i just want to show the circle which the 3 points of triangle would fall on it

OpenStudy (anonymous):

and yes cos B =-3/5

terenzreignz (terenzreignz):

Why does there have to be a circle, though? Was it included in the question?

terenzreignz (terenzreignz):

It usually helps if we make a list, or an illustration of what we know.

OpenStudy (anonymous):

sorry, it's just a condition do you call that circumcircle?

terenzreignz (terenzreignz):

Yeah, seems like it :)

OpenStudy (anonymous):

cuz the question only has words description, it gives no pic

OpenStudy (anonymous):

oh ok then it's just one of the conditions, saying the circumcircle's radius=13/2

terenzreignz (terenzreignz):

Ok, studying about circumcircles now... :P

terenzreignz (terenzreignz):

I'm hoping, sin B = 4/5. :)

OpenStudy (anonymous):

i think so , cuz cosB =-3/5 so sinB must be 4/5

terenzreignz (terenzreignz):

Yep. Well, as a result of the law of sines, \[\huge \frac{\sin B}{AC}=\frac{\sin C}{AB}\]

terenzreignz (terenzreignz):

\[\huge = \frac{\sin A}{BC}\]

terenzreignz (terenzreignz):

And it so happens these are all equal to the diameter of the circumcircle, which is twice 13/2, which is just 13.

OpenStudy (anonymous):

sorry, yeah i know that law sinB/b=13 so b=4/65 sinC/c=13,c=5 so sinC= 65 but what can i do with sinA thank you

terenzreignz (terenzreignz):

Thinking... this is making me rather dizzy :D

terenzreignz (terenzreignz):

If you can just get sin C, it would be easy :D

terenzreignz (terenzreignz):

Okay, I'm going to be bloody with this \[\large \frac{AC}{\sin B}=13 \ \ \ \ \ \ ; \ \ \ \ \ \frac{4}{5}\times 13 = AC\]

terenzreignz (terenzreignz):

Since sin B is 4/5

terenzreignz (terenzreignz):

\[\large AC = \frac{52}{5}\]

OpenStudy (anonymous):

ah i got a way, (c^2+a^2-b^2)/2ac=cosB then i can get a then cos formula again i can get cosA then i can get sinA do you think this is correct?

OpenStudy (anonymous):

sorry i didn't notice above, sorry for making you feel bloody hell

terenzreignz (terenzreignz):

\[\large \frac{\sin B}{AC}=\frac{\frac{4}{5}}{\frac{52}{5}}=\frac{4}{52}=\frac{\sin C}{AB}\] The pieces are all coming together. @ranxu6j3 I honestly have no idea, this is the first time I worked with triangles with circumcircles :D There's no need to apologise; this is fun :)

OpenStudy (anonymous):

in your calculation, we are still lack of sinA This was exactly what i was stuck for?

OpenStudy (anonymous):

@terenzreignz thanks for discussing with me

terenzreignz (terenzreignz):

Relax :) We know AB = 5, don't we? :P \[\large \frac{4}{52}=\frac{\sin C}{AB}=\frac{\sin C}{5}\] \[\large \frac{4 \times 5}{52}=\frac{5}{13}=\sin C \]

terenzreignz (terenzreignz):

And everything is coming together :D

OpenStudy (anonymous):

sorry, the question wants sinA

terenzreignz (terenzreignz):

I know. Be patient... and work on it :D \[\large \frac{\sin C}{AB}=\frac{\sin A}{BC}\]

terenzreignz (terenzreignz):

Might not have been the best way to go about this :D Your approach may be better at this point :)

OpenStudy (anonymous):

yeah, the bloody thing is that it doesn't give us BC THANKS FOR HELPING , A LOT

terenzreignz (terenzreignz):

Hang on, we can do this :)

OpenStudy (anonymous):

sorry for having to be offline now